Assume that the assumptions of the CAPM hold. The expected return and the standard deviation of the market portfolio are 7% and 14%, respectively. There are two individual stocks A and B:

Mean Return A: 4% Standard Deviation A: 18%

Mean Return B: 12% Standard Deviation B: 36%

Stock A has a correlation of 0.2 with the market portfolio.

A.What is the beta of stock A?

B.What is the risk free rate?

C.What is the beta of a portfolio with 40% in stock A and 60% in stock B?

D.What is the idiosyncratic volatility of the portfolio from part C?
Answer #1

A.

Beta = Covariance(Stock returns, market portfolio returns)/Variance(Stock Returns)

Covariance(Stock returns, market portfolio returns) = Correlation(Stock returns, market portfolio returns)*Standard Deviation of Stock Returns*Standard Deviation of Market Portfolio

B7 : X fi =B6/B4 A 1 B Stock A 4% 18% 3.24% C Stock B 12% 36% 12.96% D E Market Portfolio 7% 14% 1.96% 2 Expected Return 3 St

B7 - X fx =B6/B4 А B 2 Expected Return 3 Standard Deviation 4. Variance Stock A 0.04 0.18 =B3^2 Stock B 0.12 0.36 =C312 Marke

b.

Using CAPM

Expected Stock Return = Risk Free Rate +Beta*(Market Returns - Risk Free Rate)

Risk free rate = [Expected Stock return - Beta*Market Return]/(1 - Beta)

B9 - X V fx =(B2-B7*D2)/(1-B7) B Stock A 4% 18% 3.24% C Stock B 12% 36% 12.96% D E Market Portfolio 7% 14% 1.96% 0.20 2 Expec

B9 - x fc =(B2-B7*D2)/(1-B7) А. B. 1 2 Expected Return 3 Standard Deviation 4 Variance Stock A 0.04 0.18 =B3^2 Stock B 0.12 0

c.

Using CAPM

Beta = (Expected Return - Risk Free Rate)/(Market Returns - Risk Free Rate)

Portfolio Beta = weightof stock * betaof stock

C15 - x = fx =SUMPRODUCT(B13:B14,013:C14) F G 1 B Stock A 4% 18% 3.24% C Stock B 12% 36% 12.96% D E Market Portfolio 7% 14% 1

C15 - : X V fc =SUMPRODUCT(B13:B14,C13:C14) А B 1 2 Expected Return 3 Standard Deviation 4 Variance Stock A 0.04 0.18 =B3^2 S

D.

Idiosyncratic Volatility = Total Variance - Market Variance

B17 : x fi =D3-B16 А E р Market Portfolio 1 в в Stock A 4% 18% 3.24% 2 Expected Return 3 Standard Deviation 4 Variance с Stoc

B17 - X V fx =D3-B16 2 Expected Return 3 Standard Deviation 4. Variance Stock A 0.04 0.18 =B312 Stock B 0.12 0.36 =C312 Marke